When the first proof of 392 pages of a book of 1200 pages were 

read , the distribution of printing mistakes were found to be as 

follows

No of mistakes in a page (x) 0 1 2 3 4 5

No of pages (f) 275 72 30 7 5 2

Fit a Poisson distribution to the above data and test the goodness of



1
Expert's answer
2021-01-13T14:49:20-0500

The PDF for the poisson distribution is


"f(x)=\frac{e^{-\lambda}\lambda^x}{x!}"

Where "\lambda" is the mean. We equate the sample and population moments for the mean to obtain the parameter for the distribution



"\lambda= E(X)"


"E(X)= \frac{1}{\sum frequency}\sum x * frequency""\sum frequency =275+72+30+7+5+2 =391"

Therefore


"E(X)=\frac{1}{391}[0*275+1*72+2*30+3*7+4*5+5*2]"

"=\frac{183}{391} =0.4680"

Since "E(X)=\lambda=0.4680"


The PDF is therefore


"f(x)=\frac{e^{-0.468}0.468^x}{x!}"

Given the 391 pages, the expected frequency for the number of errors based on the poisson distribution is


"X=0: 391* \frac{e^{-0.468}0.468^0}{0!} = 244.865"

"X=1: 391* \frac{e^{-0.468}0.468^1}{1!} = 114.597"

"X=2: 391* \frac{e^{-0.468}0.468^2}{2!} = 26.816"

"X=3: 391* \frac{e^{-0.468}0.468^3}{3!} = 4.183"

"X=4: 391* \frac{e^{-0.468}0.468^4}{4!} = 0.489"

"X=5: 391* \frac{e^{-0.468}0.468^5}{5!} = 0.046"

We use the chi square test for the goodness of fit



"X^2=\sum\frac{(observed - Expected)^2}{Expected}"

"=\frac{(275-244.865)^2}{244.865} + \frac{(72-114.597)^2}{114.597} + \frac{(30-26.816)^2}{26.816} + \frac{(7-4.183)^2}{4.183} + \frac{(5-0.489)^2}{0.489} + \frac{(2-0.046)^2}{0.046}"




"X^2=146.434"


Since the sample size is greater than 200,(i.e. 391 pages were evaluated) the chi square critical value given a level of significance of 0.05 is


"X^2 _{critical\ value}=233.994"


We conclude that the model is a good fit since the calculated "X^2: 146.434 < X^2 _{critical}"